14-3A on LINK TO TEXT INTERACTIVE Review Score Review Results by Study Objective X Your answer is incorrect. Try again. Determine the missing amounts. Total Work Work Cost of Goods Manufacturing in Process in Process CostsJanuary 1) (December 31) Manufactured $121,000 $83,600 $297,000 $99,600 $323.700 $313,000$465.000 $716.000 Click if you would like to Show Work for this question: 9pn Shou Wods LINK TO TEXT Type here to searcih
Answer #1

Q1)

The formula of “cost of goods manufactured” is as below:

Cost of goods manufactured = total manufacturing cost + beginning WIP – ending WIP

                                                = 0 + 121,000 – 83,600

                                                = $37,400 (Answer)

Q2)

Beginning WIP should come by the use of “cost of goods manufactured” formula.

Cost of goods manufactured = total manufacturing cost + beginning WIP – ending WIP

323,700 = 297,000 + beginning WIP – 99,600

Beginning WIP = 323,700 – 297,000 + 99,600

                        = $126,300 (Answer)

Q3)

Ending WIP should come by the use of “cost of goods manufactured” formula.

Cost of goods manufactured = total manufacturing cost + beginning WIP – ending WIP

716,000 = 313,000 + 466,000 – ending WIP

Ending WIP = 313,000 + 466,000 – 716,000

                        = $63,000 (Answer)